OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (amistre64):

parallel lines have same homogenous parts and differ by the particular part

OpenStudy (amistre64):

keep the x and lose the constant; then calibrate a new constant with the given point

OpenStudy (amistre64):

y = x + 7 y = x + c ; use(6,-2) -2 = 6 + c -8 = c y = x-8

OpenStudy (anonymous):

y=ax+b if lines are parallel then they have the same ``a`` so y=x+7 a=1 y_2=1x+b if a point on a line then (6,-2)=(x,y) -2=1*6+b b=-8 y_2=x-8